OpenStudy (sparklestaraa):

Compare and Contrast: Below are two equations. Solve each equation and compare the two solutions. Choose the statement that is true about each solution. Equation #1 Equation #2 6 - x = 5x + 30 4x - 8 = 3x + 2 The solution to equation #1 is smaller than the solution to equation #2. The solution to equation #1 is larger than the solution to equation #2. The solution to equation #1 is the same as the solution to equation #2. None of the statements above describe solutions to equations shown.

OpenStudy (danjs):

Take it one at a time and solve for x = ...

OpenStudy (danjs):

6 - x = 5x + 30 Want to move all numbers to one side and all x's to the other side...

OpenStudy (danjs):

Ill do the first one for you, 6 - x (+x) = 5x (+x) + 30 6 = 6x + 30 (subtract 30 on both sides) 6-30 = 6x 6x = -24 (Divide both sides by 6) x = -24/6 = -4

OpenStudy (danjs):

Try doing the second one 4x - 8 = 3x + 2 What first?

OpenStudy (danjs):

take away 3x from both ends.. what do you get?

OpenStudy (sparklestaraa):

:D yes yes i do thanks so much :D